suppose $14000 is invested in 2 different funds. Part of it at 4% apr and the remaining portion at 6%apr. the total annual interest earned on the two investments is $720 Find the amount invested at 4% and at 6%?
Answer by checkley77(12844)

.06X+.04(14,000-X)=720
.06X+560-.04X=720
.02X=720-560
.02X=160
X=160/.02
X=8,000 AMOUNT INVESTED @ 6%.
14,000-8,000=6,000 AMOUNT INVESTED @ 4%
PROOF:
.06*8,000+.04*6,000=720
480+240=720
720=720